OpenStudy (anonymous):

Below you are given a function f(x) and its first and second derivatives. Use this information to solve the following.

OpenStudy (anonymous):

what is the function?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Determine the intervals where the function is concave up and concave down.

OpenStudy (anonymous):

\[f(x) = (x^2 - 4)/(x^2 +1)\]\[f'(x) = 10x/(x^2 +1)^2\]\[f''(x) = 10(1-3x^2)/(x^2 +1)^3\]

OpenStudy (anonymous):

for the first one the lowest point is at y=-4

OpenStudy (anonymous):

at f(0) i got that, isn't there a rule having to do with the second derivative i can use?

OpenStudy (anonymous):

Are you thinking of Leibniz's notation?

OpenStudy (agent0smith):

Use the second derivative to find concavity. It's concave down where f'' is negative, ie f'' < 0 Concave up where f'' is positive, f'' > 0 Find where f'' = 0 firstly.

OpenStudy (anonymous):

f' is equal to 0 when x is 2

OpenStudy (anonymous):

@Jgeurts

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

i got negative concavity for both sides of zero in [-2,3]

OpenStudy (anonymous):

so does that mean it looks something like this?